### v2

# A Geometry-Grounded Data Perimeter in Azure

Abstract
While data perimeter is ubiquitous in cybersecurity speak, it rarely defines how boundary points are arranged. In this paper we show how Azure’s blast radius ultrametric provides the distance, and how solving the Traveling Salesman Problem in this ultrametric space provides the ordering, yielding a true geometric contour: an actionable perimeter measure for SPN data plane operations de-risking.
1 Introduction
In cybersecurity and data privacy management, the notion of a data perimeter is rarely tied to a concrete geometric construction, because perimeter implies a distance function and a circular ordering of the data, both of which are usually either missing or don’t make sense for abstract datasets. To define a meaningful data perimeter in Azure, we leverage the blast radius ultrametric[7] to propose a geometry-grounded perimeter which is both efficiently computable and minimal.
We show how our data perimeter acts as a secondary sort key after blast radius stratification, to provide fine-grained risk prioritization of SPNs.
2 Background
2.1 Related Work
2.1.1 Attack Surface Model
Manadhata and Wing’s attack surface[1] pioneered the formalization of ”entry/exit points” and ”channels” as quantitative risk contributors. While their model provides a useful abstraction for OS-level systems, it is less suited to the structured permission sets typical of cloud platforms IAM like Azure RBAC.
In contrast, our approach treats the action set geometry directly, producing shape-aware summaries rather than quantitative counts.
2.1.2 Ultrametrics and Hierarchical Models
Ultrametric spaces arise naturally in settings when items are grouped by resource hierarchy. Such metrics obey the strong triangle inequality and induce a natural tree structure. Previous works on hierarchical clustering have used ultrametrics to model efficient machine learning algorithms[2], phylogenetic trees[3], and linguistic evolution[4].
Our use of ultrametrics follows this tradition, but we leverage their properties for efficient tour construction, not just visualization or embedding.
2.1.3 Seriation and metric TSP
Seriation, the process of ordering elements to reveal latent structure, has applications in archaeology, heatmap visualization, and gene expression analysis. In computational geometry, finding an optimal ordering that minimizes some distance (e.g., via the Traveling Salesman Problem) has been proposed as a generic sorting principle[5].
We contribute a novel application of this idea to the Cloud context, in Azure RBAC, showing that TSP solutions over ultrametrics are not only efficient to compute but also yield meaningful secondary metrics.
2.1.4 Perimeter-Like Measures in Security
Few existing metrics capture the notion of a perimeter over privileges. Graph-based approaches like privilege escalation trees (e.g., BloodHound[6]) focus on path enumeration and control flow, not data-action layout. Similarly, role-mining algorithms often compress permissions but ignore spatial structure.
Our notion of a data perimeter is grounded in metric space geometry, enabling fine-grained differentiation within otherwise identical blast-radius bands.
2.2 SPNs and Data Actions in Azure
Service Principals in Azure hold sets of fine-grained data actions (e.g., ReadBlob, WriteSecret) scoped to individual resources or to resource containers (resource groups, subscriptions, management groups). Ultimately, this design means that Azure Data perimeter is not a global, Tenant-wide metric: it depends on each SPN.
SPN permissions can be assigned either directly, or via groups membership. Excessive read permission may lead to a data leakage, whereas excessive write permissions may lead to data forgery. The concentration of read and write permissions across a large set of resources or resource containers under the same SPN makes risks assessment related to lateral motion particularly challenging.
2.3 Blast Radius Ultrametric
The Blast radius was introduced to address the challenge of measuring lateral motion, data leakage and data forgery of individual SPNs in Azure. It is fully impplemented in Azure Silhouette[8].
Define an ultrametric distance $d(a,b)=\tfrac{\mathrm{impact}(a,b)}{2^{2D(a,b)+1}}$, where $D$ is the Least Common Ancestor of $a$ and $b$ in Azure’s native clustering hierarchy, and $impact$ is a parameter depending on data actions[7].
As with all ultrametric distances, it satisfies the strong triangle inequality:
$\forall x,y,z\in X,\quad d(x,z)\leq\max\{d(x,y),d(y,z)\}$
The blast radius of an SPN is $B=\max_{i,j}d(x_{i},x_{j})$. It ranges from 0.0 (no permissions) to 1.0 (Tenant-wide permissions).
3 Problem Statement
The Blast Radius assigned to each SPN is a sort key to quickly identify the most risky SPNs. Due to the nature of ultrametry and the use of the $impact$ parameter, SPNs are grouped into up to 22 bands of identical Blast radius values:
  • 2 bands for Tenant-wide radius
  • 12 bands for Management groups
  • 2 bands for subscriptions
  • 2 bands for resource groups
  • 2 bands for resources
  • 2 bands for resource parts
The Blast radius stratifies SPNs into coarse bands, so many SPNs tie on the same bands, leaving no intra-band ranking.
We seek a secondary sorting key that leverages the ultrametric structure to distinguish SPNs further.
4 Data Perimeter
4.1 Formulation
Given data actions ${x_{1},\dots,x_{n}}$ with ultrametric $d$, the Data Perimeter is defined as the minimal circular path covering all actions under ultrametric distance.
$P=\min_{\pi\in S_{n}}\sum_{i=1}^{n}d(x_{\pi(i)},x_{\pi(i+1)}),\quad\text{with }x_{\pi(n+1)}=x_{\pi(1)}$
Where $\pi$ is a permutation of the data action indices.
This amounts to solving a metric Traveling Salesman Problem (metric TSP). A major benefit of using the ultrametric TSP formulation for computing the data perimeter is that the problem, typically NP-hard in general, becomes efficiently solvable. In ultrametric spaces, the hierarchical structure imposed by the strong triangle inequality enables polynomial-time solutions to the metric TSP. This stems from the fact that shortest tours respect the natural tree structure of the space. That is, the tour effectively follows a traversal of the leaves of the ultrametric tree with minimal backtracking.
So, the data perimeter is not only geometrically grounded but also computationally tractable in our setting. The polynomial complexity allows for practical scaling even with large sets of data actions or SPNs. It is the sum of distances in the tour, with wraparound from the last data point to the first:
$P=\sum_{i=1}^{n}d(x_{i},x_{i+1}),\quad\text{with }x_{n+1}=x_{1}$
Note that:
  • Ultrametric TSP admits multiple optimal tours, so the circular ordering required for constructing a perimeter is not unique, but all tour lengths are equal. They form an equivalence class.
  • Because of the wraparound, a tour length doesn’t depend on its starting point.
4.1.1 Ultrametric TSP pseudo-code
Input: A set of data actions $A=\{a_{1},\dots,a_{n}\}$ with distance function $d(a,b)$
Output: Perimeter length $L$
  1. 1.
    Let $P\leftarrow[\,]$ (Tour list)
  2. 2.
    Find $a_{\text{start}}\in A$ such that
    $a_{\text{start}}=\arg\max_{a\in A}\max_{b\in A}d(a,b)$
  3. 3.
    Append $a_{\text{start}}$ to $P$
  4. 4.
    Let $U\leftarrow A\setminus\{a_{\text{start}}\}$ (Unvisited actions)
  5. 5.
    while $U\neq\emptyset$ do
    1. (a)
      Let $a_{\text{current}}\leftarrow$ last element of $P$
    2. (b)
      Find $a_{\text{next}}\in U$ such that $d(a_{\text{current}},a_{\text{next}})$ is minimized
    3. (c)
      Append $a_{\text{next}}$ to $P$
    4. (d)
      Remove $a_{\text{next}}$ from $U$
  6. 6.
    Append $a_{\text{start}}$ to $P$ to complete the cycle
  7. 7.
    Initialize $L\leftarrow 0$
  8. 8.
    for $i=1$ to $|P|-1$ do
    1. (a)
      $L\leftarrow L+d(P_{i},P_{i+1})$
  9. 9.
    return $L$
4.1.2 Limitation
Silhouette implements Blast radii calculation over the native Azure Tenant hierarchy, as well as on a family of alternate hierarchies. This option lets risk officers tailor blast radii to the local needs of their organization: they can for instance use a special alternate hierarchy to represent the state of a reorganization, a merger or acquisition, a carve-out, …
When hierarchical families are defined, Silhouette computes the Blast radius as the pointwise infimum in all hierarchies to yield a tighter, smaller Blast Radius than the native one. In doing so, it must be noted that the pointwise infimum is not ultrametric, hence the data Perimeter should be treaded with care in this situation. Falling back to the native tenant hierarchy is highly recommended.
Here is a small counter-example showing the lack of ultrametry:
Let $X=\{x,y,z\}$, and define two ultrametrics $d_{1}$ and $d_{2}$ on $X$ as follows:
$d_{1}(x,y)=2,\quad d_{1}(y,z)=1,\quad d_{1}(x,z)=2,$
$d_{2}(x,y)=1,\quad d_{2}(y,z)=2,\quad d_{2}(x,z)=2.$
Each $d_{i}$ satisfies the ultrametric inequality:
$d_{i}(a,c)\leq\max\{d_{i}(a,b),d_{i}(b,c)\}\quad\text{for all }a,b,c\in X.$
Now define $d_{3}$ as the pointwise minimum:
$d_{3}(a,b):=\min\{d_{1}(a,b),d_{2}(a,b)\}.$
This gives:
$d_{3}(x,y)=\min\{2,1\}=1,$
$d_{3}(y,z)=\min\{1,2\}=1,$
$d_{3}(x,z)=\min\{2,2\}=2.$
However, the strong triangle inequality fails:
$d_{3}(x,z)=2\not\leq\max\{d_{3}(x,y),d_{3}(y,z)\}=\max\{1,1\}=1.$
5 Mean distance
The average distance between any two data actions is the metric that springs to mind for complementing the Blast radius. It captures global dispersion but ignores local structure and chaining. In this section, we explain why this doesn’t make it a good secondary key candidate for sorting SPNs.
5.1 Comparison with Data Perimeter
To understand the benefits of the data perimeter over the mean, we sample 1285 SPNs that we sort by Blast band:
  1. 1.
    For each SPN we compute a spread_ratio defined as $\frac{\text{P}}{n\cdot\mu_{\mathrm{mean}}}$, where $n$ is the data actions count held by this SPN. The reason why we divide $P$ by $n$ is to make it comparable with $\mu_{\mathrm{mean}}$.
  2. 2.
    We group SPNs by band
  3. 3.
    We calculate the average spread ratio in each band.
Table 1 summarizes our findings. Blast radii are first shuffled, then anonymized into Band IDs.
Band ID SPNs count Spread Ratio Regime
I 33 1.00000 Tight
II 69 0.998993 Tight
III 4 0.820245 Dispersed
IV 2 1.00000 Tight
V 110 0.99186 Tight
VI 8 0.978792 Dispersed
VII 228 1.00000 Tight
VIII 19 1.00000 Tight
IX 3 1.00000 Tight
X 505 0.997459 Tight
XI 6 1.00000 Tight
XII 70 0.987850 Tight
XIII 2 1.00000 Tight
XIV 16 1.00000 Tight
XV 28 0.581303 Dispersed
XVI 66 0.806384 Dispersed
XVII 2 1.00000 Tight
XVIII 75 0.950056 Dispersed
XIX 5 0.956395 Dispersed
XX 2 1.00000 Dispersed
XXI 32 0.888986 Dispersed
Table 1: In-band spread ratios (data perimeter vs. mean) per shuffled and anonymized blast radius band.
5.1.1 Discussion
The in-band spread_ratio values (Table 1) fall into two distinct regimes:
Tightly clustered SPNs (blast radius $\lesssim 10^{-4}$)
In bands VII-XIV, and other small radius groups, the normalized data perimeter and the mean coincide (spread_ratio $\approx 1.0$). Here, each SPN’s permission set is nearly point-like, so any reasonable ordering of actions (whether implicit in the mean or made explicit via TSP) produces the same minimal contour length.
Dispersed SPNs (blast radius $\gtrsim 10^{-4}$)
As blast radii grow, we see a systematic divergence (spread_ratio $<1$). For example:
  • Band XV: spread_ratio 0.58
  • Band XVI: spread_ratio 0.81
  • Band XVIII: spread_ratio 0.95
This reflects the fundamental difference between our two metrics.
  • The mean sums large, disjoint jumps and thus overestimates overall spread.
  • The TSP-based perimeter constructs a continuous tour, chaining cluster centers and “shortcutting” around outliers. By exploiting the strong triangle inequality, it compresses lengthy detours into shorter, hierarchical paths.
Implications for SPN prioritization
When many SPNs share the same blast radius, the data perimeter provides a fine-grained, geometry-aware tiebreaker. It surfaces principals whose permission sets exhibit branching or elongated structures (profiles that pose higher risk in practice) whereas the mean alone would treat them identically to more compact SPNs.
This richer signal enables security teams to target reviews and policy adjustments not just based on “how far” an SPN can reach, but on “how” its permissions are organized. It confirms that TSP-derived perimeter is not just another average: it captures the global connectivity structure of data actions.
6 Conclusion
We introduced the data perimeter as a geometry-grounded refinement of the blast radius, capturing the spatial layout and redundancy of data actions assigned to service principals (SPNs). Unlike scalar averages such as the ultrametric mean, the perimeter is sensitive to seriation, clustering, and the underlying topology of permission sets.
By leveraging ultrametric distance and a tractable TSP solver, we demonstrated that the data perimeter can be computed efficiently and yields a non-arbitrary circular ordering of actions. This ordering is meaningful, not merely heuristic: it reflects a class of equivalent perimeters that respect the hierarchical structure of permissions, compressing global dispersion into a single tour.
Our small comparative analysis across blast radius bands revealed that while ultrametric mean and perimeter agree in tightly scoped SPNs, they diverge significantly as permission sets grow complex. The data perimeter remains stable and informative even in the presence of branching or fragmented access patterns, highlighting its suitability as a secondary sorting key for SPNs sharing similar blast magnitude.
This work bridges abstract privilege metrics with actionable geometric intuition. It enables analysts to go beyond “how much power” toward “how this power is organized,” providing a new lens to prioritize, visualize, and reduce risk in large-scale cloud environments.

### v3

@@
 While Data Perimeter is ubiquitous in cybersecurity speak, it rarely defines how boundary points are arranged.
-In this paper we show how Azure’s blast radius ultrametric provides the distance, and how solving the Traveling Salesman Problem in this ultrametric space provides the ordering, yielding a true geometric contour: an actionable perimeter measure for SPN data plane operations de-risking.
+In this paper we show how Azure’s blast radius ultrametric provides the distance, and how solving the Traveling Salesman Problem in this ultrametric space provides the ordering, yielding a true geometric contour: an actionable perimeter measure for Service Principals (SPNs) data plane operations de-escalation.
 1 Introduction
@@
 Previous works on hierarchical clustering have used ultrametrics to model efficient machine learning algorithms[2], phylogenetic trees[3], and linguistic evolution[4].
-Our use of ultrametrics follows this tradition, but we leverage their properties for efficient tour construction, not just visualization or embedding.
+Our use of ultrametrics follows this tradition, but we leverage their properties for efficient tour construction (not just visualization).
 2.1.3 Seriation and metric TSP
@@
 2.3 Blast Radius Ultrametric
-The Blast radius was introduced to address the challenge of measuring lateral motion, data leakage and data forgery of individual SPNs in Azure. It is fully impplemented in Azure Silhouette[8].
+The Blast radius was introduced to address the challenge of measuring lateral motion, data leakage and data forgery of individual SPNs in Azure. It is fully implemented in Azure Silhouette[8].
 Define an ultrametric distance $d(a,b)=\tfrac{\mathrm{impact}(a,b)}{2^{2D(a,b)+1}}$, where $D$ is the Least Common Ancestor of $a$ and $b$ in Azure’s native clustering hierarchy, and $impact$ is a parameter depending on data actions[7].
@@
 $\forall x,y,z\in X,\quad d(x,z)\leq\max\{d(x,y),d(y,z)\}$
-The blast radius of an SPN is $B=\max_{i,j}d(x_{i},x_{j})$. It ranges from 0.0 (no permissions) to 1.0 (Tenant-wide permissions).
+The blast radius of an SPN is the diameter of $d()$: $\max_{i,j}d(x_{i},x_{j})$. It ranges from 0.0 (no permissions) to 1.0 (Tenant-wide permissions).
 3 Problem Statement
@@
 That is, the tour effectively follows a traversal of the leaves of the ultrametric tree with minimal backtracking.
-So, the data perimeter is not only geometrically grounded but also computationally tractable in our setting.
-The polynomial complexity allows for practical scaling even with large sets of data actions or SPNs.
-It is the sum of distances in the tour, with wraparound from the last data point to the first:
-$P=\sum_{i=1}^{n}d(x_{i},x_{i+1}),\quad\text{with }x_{n+1}=x_{1}$
-Note that:
-  • Ultrametric TSP admits multiple optimal tours, so the circular ordering required for constructing a perimeter is not unique, but all tour lengths are equal. They form an equivalence class.
-  • Because of the wraparound, a tour length doesn’t depend on its starting point.
+As will be shown in the ultrametric TSP pseudo-code below, the Data Perimeter is not only geometrically grounded but also computationally tractable in our setting.
 4.1.1 Ultrametric TSP pseudo-code
@@
   2. 2.
-    Find $a_{\text{start}}\in A$ such that
-    $a_{\text{start}}=\arg\max_{a\in A}\max_{b\in A}d(a,b)$
+    Pick $a_{\text{start}}\in A$ at random
   3. 3.
@@
     return $L$
-4.1.2 Limitation
+4.1.2 Notes
+  • Ultrametric TSP admits multiple optimal tours, so the circular ordering required for constructing a perimeter is not unique, but all tour lengths are equal. They form an equivalence class.
+  • Because of the wraparound, a tour length doesn’t depend on its starting point.
+4.1.3 Known Limitations
 Silhouette implements Blast radii calculation over the native Azure Tenant hierarchy, as well as on a family of alternate hierarchies.
@@
 When hierarchical families are defined, Silhouette computes the Blast radius as the pointwise infimum in all hierarchies to yield a tighter, smaller Blast Radius than the native one.
-In doing so, it must be noted that the pointwise infimum is not ultrametric, hence the data Perimeter should be treaded with care in this situation.
+In doing so, it must be noted that the pointwise infimum is not ultrametric, hence the data Perimeter should be treated with care in this situation.
 Falling back to the native tenant hierarchy is highly recommended.
@@
 $d_{3}(x,z)=2\not\leq\max\{d_{3}(x,y),d_{3}(y,z)\}=\max\{1,1\}=1.$
-5 Mean distance
+5 Mean distance ($\mu_{\mathrm{mean}}$)
 The average distance between any two data actions is the metric that springs to mind for complementing the Blast radius.
-It captures global dispersion but ignores local structure and chaining.
+Unfortunately, if it captures global dispersion, it ignores local structure and chaining.
 In this section, we explain why this doesn’t make it a good secondary key candidate for sorting SPNs.
@@
     For each SPN we compute a spread_ratio defined as $\frac{\text{P}}{n\cdot\mu_{\mathrm{mean}}}$, where $n$ is the data actions count held by this SPN.
-The reason why we divide $P$ by $n$ is to make it comparable with $\mu_{\mathrm{mean}}$.
+The perimeter $P$ is normalized by the action count $n$ to maintain dimensional consistency with $\mu_{\mathrm{mean}}$.
   2. 2.
@@
 When many SPNs share the same blast radius, the Data Perimeter provides a fine-grained, geometry-aware tiebreaker.
-It surfaces principals whose permission sets exhibit branching or elongated structures (profiles that pose higher risk in practice) whereas the mean alone would treat them identically to more compact SPNs.
+It surfaces principals whose permission sets exhibit branching or elongated ultrametric structures whereas the mean alone would treat them identically to more compact SPNs.
 This richer signal enables security teams to target reviews and policy adjustments not just based on “how far” an SPN can reach, but on “how” its permissions are organized.
@@
 It enables analysts to go beyond “how much power” toward “how this power is organized,” providing a new lens to prioritize, visualize, and reduce risk in large-scale cloud environments.
+Appendix: Minimal Data Perimeter
+In this work, we have argued that the Data Perimeter in Azure security models behaves not just as a metaphor, but as a measurable geometric quantity.
+To further give substance to this geometric view, we identify the shapes corresponding to the smallest Data Perimeter.
+Ultracycle and Minimal Perimeter
+Given an ultrametric space $(X,d)$ with $n$ points, an ultracycle is a configuration where all pairwise distances are equal: $d(x_{i},x_{j})=\delta$ for all distinct $i,j\in\{1,2,\ldots,n\}$ and some constant $\delta>0$.
+In any ultrametric space, ultracycles uniquely minimize the TSP tour length among all configurations with the same number of points.
+Proof
+Let $(X,d)$ be an ultrametric space and consider $n$ points $\{x_{1},x_{2},\ldots,x_{n}\}\subset X$.
+Case 1: Ultracycle Configuration
+Suppose all pairwise distances are equal: $d(x_{i},x_{j})=\delta$ for all $i\neq j$.
+For any TSP tour $\pi$ visiting all points exactly once, the tour length is:
+$P_{\text{ultra}}=\sum_{i=1}^{n}d(x_{\pi(i)},x_{\pi(i+1)})=\sum_{i=1}^{n}\delta=n\delta$ (1)
+where $x_{\pi(n+1)}:=x_{\pi(1)}$.
+Since all edges have the same weight $\delta$, every possible tour has identical length $n\delta$.
+Case 2: Non-Ultracycle Configuration
+Now suppose the configuration is not an ultracycle. Then there exist distinct indices $i,j,k,\ell$ such that:
+$d(x_{i},x_{j})\neq d(x_{k},x_{\ell})$ (2)
+Let $d_{\min}=\min_{i\neq j}d(x_{i},x_{j})$ and $d_{\max}=\max_{i\neq j}d(x_{i},x_{j})$.
+By assumption, $d_{\min}<d_{\max}$.
+Lemma: in any ultrametric space, if $d(a,b)<d(a,c)$, then $d(b,c)=d(a,c)$.
+Proof of Lemma: By the strong triangle inequality: $d(a,c)\leq\max\{d(a,b),d(b,c)\}$.
+So we must have $d(a,c)\leq d(b,c)$.
+Similarly, $d(b,c)\leq\max\{d(b,a),d(a,c)\}$. But $d(a,b)<d(a,c)$, so $d(b,c)\leq d(a,c)$.
+Therefore, $d(b,c)=d(a,c)$. This proves the Lemma.
+Now, let $S_{\min}=\{(i,j):d(x_{i},x_{j})=d_{\min}\}$ be the set of point pairs with minimum distance.
+Claim: Any TSP tour must include at least one edge with distance $>d_{\min}$.
+Proof of Claim: Suppose, for contradiction, that there exists a tour using only edges of distance $d_{\min}$.
+This would mean we can connect all $n$ points using only edges from $S_{\min}$, forming a Hamiltonian cycle.
+However, by the ultrametric property (Lemma), if $d(x_{i},x_{j})=d_{\min}$ and $d(x_{j},x_{k})=d_{\min}$, then either:
+  • $d(x_{i},x_{k})=d_{\min}$ (all three points are equidistant), or
+  • $d(x_{i},x_{k})>d_{\min}$ (the points form a non-equidistant triple)
+If all points were connected by minimum-distance edges, then by transitivity of the equidistance relation in ultrametric spaces, all pairwise distances would equal $d_{\min}$, contradicting our assumption that the configuration is not an ultracycle.
+Therefore, any Hamiltonian cycle must include at least one edge with distance $>d_{\min}$.
+Completing the Proof:
+Let $P_{\text{non-ultra}}$ be the length of an optimal TSP tour in the non-ultracycle configuration.
+Since the tour must include at least one edge of length $>d_{\min}$, and the remaining $(n-1)$ edges have length $\geq d_{\min}$:
+$P_{\text{non-ultra}}>d_{\min}+(n-1)\cdot d_{\min}=n\cdot d_{\min}$ (3)
+For an ultracycle with the same minimum distance $d_{\min}$ (i.e., $\delta=d_{\min}$):
+$P_{\text{ultra}}=n\cdot d_{\min}$ (4)
+Therefore: $P_{\text{non-ultra}}>P_{\text{ultra}}$.
+This proves that ultracycles uniquely minimize the TSP tour length in ultrametric spaces.
+Security Implications
+In Azure’s permission model, a Principal with $n$ data actions all scoped to the same resource (forming an ultracycle at depth $d$ in the hierarchy) achieves the minimal Data Perimeter, representing optimal privilege containment.
+This geometric result aligns with the theoretical foundation for the least-privilege principle: permissions should be clustered as tightly as possible in the resource hierarchy to minimize potential lateral movement exposure.

### v4

@@
 That is, the tour effectively follows a traversal of the leaves of the ultrametric tree with minimal backtracking.
-As will be shown in the ultrametric TSP pseudo-code below, the Data Perimeter is not only geometrically grounded but also computationally tractable in our setting.
+As shown in the ultrametric TSP pseudo-code implementing the nearest-neighbor algorithm below, the Data Perimeter is not only geometrically grounded but also computationally tractable in our setting.
 4.1.1 Ultrametric TSP pseudo-code
@@
 Ultracycle and Minimal Perimeter
-Given an ultrametric space $(X,d)$ with $n$ points, an ultracycle is a configuration where all pairwise distances are equal: $d(x_{i},x_{j})=\delta$ for all distinct $i,j\in\{1,2,\ldots,n\}$ and some constant $\delta>0$.
+Given an ultrametric space $(X,d)$ with $n$ points, an ultracycle is a configuration where all pairwise distances are equal: $d(x_{i},x_{j})=\xi$ for all distinct $i,j\in\{1,2,\ldots,n\}$ and some constant $\xi>0$.
 In any ultrametric space, ultracycles uniquely minimize the TSP tour length among all configurations with the same number of points.
@@
 Case 1: Ultracycle Configuration
-Suppose all pairwise distances are equal: $d(x_{i},x_{j})=\delta$ for all $i\neq j$.
+Suppose all pairwise distances are equal: $d(x_{i},x_{j})=\xi$ for all $i\neq j$.
 For any TSP tour $\pi$ visiting all points exactly once, the tour length is:
-$P_{\text{ultra}}=\sum_{i=1}^{n}d(x_{\pi(i)},x_{\pi(i+1)})=\sum_{i=1}^{n}\delta=n\delta$ (1)
-where $x_{\pi(n+1)}:=x_{\pi(1)}$.
-Since all edges have the same weight $\delta$, every possible tour has identical length $n\delta$.
+$P_{\text{ultra}}=\sum_{i=1}^{n}d(x_{\pi(i)},x_{\pi(i+1)})=\sum_{i=1}^{n}\xi=n\xi,\text{ where }x_{\pi(n+1)}:=x_{\pi(1)}.$
+Every possible tour has identical length $n\xi$.
 Case 2: Non-Ultracycle Configuration
 Now suppose the configuration is not an ultracycle. Then there exist distinct indices $i,j,k,\ell$ such that:
-$d(x_{i},x_{j})\neq d(x_{k},x_{\ell})$ (2)
-Let $d_{\min}=\min_{i\neq j}d(x_{i},x_{j})$ and $d_{\max}=\max_{i\neq j}d(x_{i},x_{j})$.
-By assumption, $d_{\min}<d_{\max}$.
-Lemma: in any ultrametric space, if $d(a,b)<d(a,c)$, then $d(b,c)=d(a,c)$.
-Proof of Lemma: By the strong triangle inequality: $d(a,c)\leq\max\{d(a,b),d(b,c)\}$.
-So we must have $d(a,c)\leq d(b,c)$.
-Similarly, $d(b,c)\leq\max\{d(b,a),d(a,c)\}$. But $d(a,b)<d(a,c)$, so $d(b,c)\leq d(a,c)$.
-Therefore, $d(b,c)=d(a,c)$. This proves the Lemma.
-Now, let $S_{\min}=\{(i,j):d(x_{i},x_{j})=d_{\min}\}$ be the set of point pairs with minimum distance.
+$d(x_{i},x_{j})\neq d(x_{k},x_{\ell})$
+Let $d_{\min}=\min_{i\neq j}d(x_{i},x_{j})$.
 Claim: Any TSP tour must include at least one edge with distance $>d_{\min}$.
-Proof of Claim: Suppose, for contradiction, that there exists a tour using only edges of distance $d_{\min}$.
-This would mean we can connect all $n$ points using only edges from $S_{\min}$, forming a Hamiltonian cycle.
-However, by the ultrametric property (Lemma), if $d(x_{i},x_{j})=d_{\min}$ and $d(x_{j},x_{k})=d_{\min}$, then either:
-  • $d(x_{i},x_{k})=d_{\min}$ (all three points are equidistant), or
-  • $d(x_{i},x_{k})>d_{\min}$ (the points form a non-equidistant triple)
-If all points were connected by minimum-distance edges, then by transitivity of the equidistance relation in ultrametric spaces, all pairwise distances would equal $d_{\min}$, contradicting our assumption that the configuration is not an ultracycle.
-Therefore, any Hamiltonian cycle must include at least one edge with distance $>d_{\min}$.
-Completing the Proof:
-Let $P_{\text{non-ultra}}$ be the length of an optimal TSP tour in the non-ultracycle configuration.
-Since the tour must include at least one edge of length $>d_{\min}$, and the remaining $(n-1)$ edges have length $\geq d_{\min}$:
-$P_{\text{non-ultra}}>d_{\min}+(n-1)\cdot d_{\min}=n\cdot d_{\min}$ (3)
-For an ultracycle with the same minimum distance $d_{\min}$ (i.e., $\delta=d_{\min}$):
-$P_{\text{ultra}}=n\cdot d_{\min}$ (4)
-Therefore: $P_{\text{non-ultra}}>P_{\text{ultra}}$.
-This proves that ultracycles uniquely minimize the TSP tour length in ultrametric spaces.
+Proof of Claim:
+Suppose, for contradiction, that there exists a tour using only edges of distance $d_{\min}$.
+For any three points $x_{i}$, $x_{j}$, $x_{k}$:
+  • If $d(x_{i},x_{j})=d_{min}$ and $d(x_{j},x_{k})=d_{min}$, then $d(x_{i},x_{k})\leq\max\{d(x_{i},x_{j}),d(x_{j},x_{k})\}=d_{min}$
+  • Since $d_{min}$ is the global minimum distance, $d(x_{i},x_{k})\geq d_{min}$
+  • Thus, $d(x_{i},x_{k})=d_{min}$
+By induction, any two points connected by a path of $d_{min}$ edges must be at distance $d_{min}$.
+Since the tour connects all $n$ points via $d_{min}$ edges, all pairwise distances must be $d_{min}$, forming and ultracycle.
+This contradicts the non-ultracycle assumption.
+Ultracycles uniquely minimize the TSP tour length in ultrametric spaces.
 Security Implications